The larger of two numbers is three more than five times the smaller. If the sum of the two numbers is 27, find thae larger number
Answer by RAY100(1637)

Let x = larger number,,,,and y = other number
.
x+y = 27,,,,,,y=27-x
.
x= 5y+3,,,,,,,subst
.
x = 5(27-x) +3
.
x= 135 -5x +3
.
6x = 138
.
x= 23,,,,,,,,y= 27-23 = 4
.
.
check..(23,4),,,,,23+4 =27,,,ok
.
23= 5*4 +3 =23,,,,,,ok
.